OpenStudy (anonymous):

I'm having a hard time with this conversion. 10 yd 4ft 4in =____________in?

ganeshie8 (ganeshie8):

rewrite the expression as below : 10 yd + 4 ft + 4 in. use : 1 yd = 36 in 1 ft = 12 inch

OpenStudy (anonymous):

So.... the answer is 412 inches...total??? Is that right?

ganeshie8 (ganeshie8):

correct

OpenStudy (anonymous):

Yeeeeeh....thank you

OpenStudy (anonymous):

I need to change this number 0.05 to a fractions/mixed number?

ganeshie8 (ganeshie8):

\[0.05 = 0.05 x 100/100 = 5/100 = 1/20\]